# AVTENet: Audio-Visual Transformer-based Ensemble Network Exploiting Multiple Experts for Video Deepfake Detection

### v1

@@ IV-D5 Comparison of AVTENet with two different AVNs
Our proposed AVTENet model is an ensemble of an AST-based AN model, a ViViT-based VN model, and an AV-HuBERT-based AVN model. Although $AVTENet_{ff}$ performs well on all test sets, its performances on the faceswap and fsgan test sets are relatively low, where the fake instances are only visually manipulated while audio tracks are real. We speculate that the reason may be that the AV-HuBERT feature extractor in AVN only extracts visual information from lip images and ignores information outside the lip region. Therefore, we also implement another AVN based on AST and ViViT. Here, we compare AVTENet (AST_ViViT_AV-HuBERT) and AVTENet (AST_ViViT_AST&ViViT). From Table XI, it is clear that although the AST_ViViT_AST&ViViT-based $AVTENet_{mv}$ and $AVTENet_{asf}$ models do slightly improve the performance on the faceswap test set, the AST_ViViT_AST&ViViT-based AVTENet models perform worse than the AST_ViViT_AV-HuBERT-based AVTENet models. The reason may be that since the AST and ViViT models are already used in the AN and VN models, they cannot provide additional information in the AVN model. Surprisingly, the AST_ViViT_AST&ViViT-based AVTENet models almost fail on the RTVC test set. To fully utilize face images in the AVN model, we may need to apply new pre-trained visual models.
IV-D6 Comparison of AVTENet and other Models
@@
Finally, as shown in Fig. 7 and Table XII, we conduct a thorough comparison of our proposed $AVTENet_{ff}$ model with various existing uni-modal, multi-modal, fusion, and ensemble methods. Several uni-modal, multi-modal and ensemble methods were presented in [56], but most of them failed to detect acoustic and visual manipulations in videos well. VGG16, a uni-modal approach exclusively trained on visual data, outperforms all other uni-modal approaches reported in [56] for video forgery detection with an accuracy of 0.81. Xception, another uni-modal method reported in [56], is trained exclusively on the audio modality and achieves an accuracy of only 0.76. Similarly, Lip Forensics [35] exploits visual lip movements to detect forgeries in videos with an accuracy of 0.76. Although Ensemble (soft-voting and hard-voting), Multi-modal-1, Multi-modal-2, and Multi-modal-3 utilize both audio and visual modalities for deepfake detection, their performance is still unsatisfactory. Not-made-for-each other [16] is an audio-visual dissonance-based approach that detects video forgeries with an accuracy of 0.69. The multi-modal ensemble approach proposed in [10] utilizes three CNN-based classifiers to identify forgeries and achieves an accuracy of 0.89. Another audio-visual approach proposed in [11] successfully identifies fake videos using lip synchronization, and the model achieves an accuracy of 0.94. Moreover, Multimodaltrace is an ensemble approach proposed in [58], which achieves an accuracy of 0.929. AVFakeNet [59] is also an ensemble approach using audio and video transformer encoders, which achieves an accuracy of 0.934. AVoiD-DF [60] is another audio-visual fusion approach, but it only achieves an accuracy of 0.837. MIS-AVioDD [61] is also a fusion approach that jointly uses the modality invariant and modality-specific representations to detect the audio-visual forgery and achieves an accuracy of 0.962. From Table XII, it is evident that our ensemble approach outperforms other existing models, achieving an accuracy of 0.99 on Testset-II of the FakeAVCeleb dataset, which is a new SOTA performance.
V Conclusions and Future Work
@@
In this work, we demonstrated the effectiveness of ensemble learning and studied deepfake video detection from an audio-visual perspective. We proposed a novel audio-visual transformer-based ensemble learning solution for effective and scalable video forgery detection, especially for videos containing acoustic and visual manipulations, which cannot rely on existing methods because they can only detect single-modal manipulations or lack training diversity. Specifically, we have devised a transformer-based model with a powerful detection ability due to its sequence modeling, parallelization, and attention mechanisms. We have illustrated the performances of our multi-modal transformer-based ensemble model on several manipulation techniques and compared it with various existing uni-modal, multi-modal, fusion, and ensemble models. Our model achieves state-of-the-art performance on the FakeAVCeleb dataset. It should be emphasized that the proposed solution is extensible via self-supervised learning models with large amounts of training data. To further improve upon the results presented in the paper, we plan to advance this research by integrating more recent self-supervised learning models.

### v2

@@ IV-D6 Comparison of AVTENet with Different AVNs
Our proposed AVTENet model is an ensemble of an AST-based AN model, a ViViT-based VN model, and an AV-HuBERT-based AVN model. Although $AVTENet_{ff}$ performs well on all test sets, its performances on the faceswap and fsgan test sets are relatively low, where the fake instances are only visually manipulated while audio tracks are real. We speculate that the reason may be that the AV-HuBERT feature extractor in AVN only extracts visual information from lip images and ignores information outside the lip region. Therefore, we also implement another AVN based on AST and ViViT. Here, we compare AVTENet (AST_ViViT_AV-HuBERT) and AVTENet (AST_ViViT_AST&ViViT). From Table XI, although the AST_ViViT_AST&ViViT-based $AVTENet_{mv}$ and $AVTENet_{asf}$ models do slightly improve the performance on the faceswap test set, the AST_ViViT_AST&ViViT-based AVTENet models perform worse than the AST_ViViT_AV-HuBERT-based AVTENet models. The AST and ViViT models are already used in the AN and VN models. Therefore, they cannot provide additional information on the AVN model. Surprisingly, the AST_ViViT_AST&ViViT-based AVTENet models almost fail on the RTVC test set. To fully utilize face images in the AVN model, we may need to apply new pre-trained visual models.
IV-D7 Comparison of AVTENet and Other Models
@@
Finally, as shown in Fig. 11 and Table XII, we conduct a thorough comparison of our proposed $AVTENet_{ff}$ model with various existing unimodal, multimodal, fusion, and ensemble methods. Several unimodal, multimodal and ensemble methods were presented in [69], but most of them failed to detect acoustic and visual manipulations in videos well. VGG16, a unimodal approach exclusively trained on visual data, outperforms all other unimodal approaches reported in [69] for video forgery detection with an accuracy of 0.81. Xception, another unimodal method reported in [69], is trained exclusively on the audio modality and achieves an accuracy of only 0.76. Similarly, Lip Forensics [42] exploits visual lip movements to detect forgeries in videos with an accuracy of 0.76. Although Ensemble (soft-voting and hard-voting), Multimodal-1, Multimodal-2, and Multimodal-3 utilize both audio and visual modalities for deepfake detection, their performance is still unsatisfactory. Not-made-for-each other [15] is an audio-visual dissonance-based approach that detects video forgeries with an accuracy of 0.69. The multimodal ensemble approach proposed in [9] utilizes three CNN-based classifiers to identify forgeries and achieves an accuracy of 0.89. Another audio-visual approach proposed in [10] successfully identifies fake videos using lip synchronization, and the model achieves an accuracy of 0.94. Moreover, the ensemble approach Multimodaltrace[70] achieves an accuracy of 0.93. AVFakeNet[71] is also an ensemble approach using audio and video transformer encoders, which achieves an accuracy of 0.93. AVoiD-DF [72] is another audio-visual fusion approach, but it only achieves an accuracy of 0.84. Furthermore, MIS-AVioDD[73] is also a fusion approach that jointly uses the modality invariant and modality-specific representations to detect audio-visual forgery and achieves an accuracy of 0.96. Table XII shows that our ensemble approach outperforms other existing models, with an accuracy of 0.99 on Testset-II of the FakeAVCeleb dataset, which is a new SOTA performance.
TABLE XIII: Performance comparison of the proposed model and other models tested on the DFDC test set.
Model AUC Score (%)
LipForensics [65] 53.10
CViT [74] 56.70
MesoNet [11] 41.20
MDS [15] 73.80
AVoiD-DF [72] 80.60
Proposed model ($\text{AVTENet}_{\text{ff}}$) 64.10
Proposed model ($\text{AVTENet}_{\text{ff}}$) (on the frontal face subset) 83.84
IV-D8 Comparison of AVTENet and Humans
We conducted a preliminary comparison of AVTENet’s detection capabilities with human cognitive proficiency in detecting audio-visual deepfakes using a manually curated subset of 30 videos (15 real videos and 15 fake videos) from the FakeAVCeleb test set to ensure a diverse and challenging evaluation set, with all fake videos featuring audio-visual manipulation.
The videos were presented to 24 subjects in three phases to classify each video as real or fake. In the first phase (audio-only), participants judged whether the video was authentic or not by listening to the audio track alone. In the second phase (video-only), participants judged the authenticity of the video using only visual input. Finally, in the third phase (audio-visual), participants integrated the audio and visual streams to classify the video as real or fake. As shown in Table XIV, humans rely on their own perception and reasoning capabilities to achieve an accuracy of 0.58 under the audio-only condition, 0.63 under the video-only condition, and 0.76 under the audio-visual condition. From Table XIV, it is clear that, under three conditions, AST (audio-only network), ViViT (video-only network), and AVTENet (audio-visual network) outperform humans (0.77 vs 0.58, 0.83 vs 0.63, and 0.93 vs 0.76), respectively. The results highlight the importance of multimodal sensory integration for enhancing human and model performance in video forgery detection; especially for detection models, combining audio and visual cues can lead to significant performance improvements.
TABLE XIV: Comparison of humans and models in video forgery detection.
Humans Models
Modalities
Accuracy
Modalities
Accuracy
Audio-Only 0.58 AST (Audio-Only) 0.77
Video-Only 0.63 ViViT (Video-Only) 0.83
Audio-visual 0.76 AVTENet (Audio-visual) 0.93
IV-D9 Generalization Test
Generalization remains a significant challenge for video forgery detection tasks, including multimodal deepfake detection. To evaluate the generalization ability of our proposed method, we directly test the proposed model trained on the FakeAVCeleb dataset on the DFDC test set, and compare it with other models under the same setting. We extract the test videos and preprocess them following [15] and [75]. As shown in Table XIII, our model achieves an AUC (area under the receiver operating characteristic curve) score of 64.10%, which is better than the performance of LipForensics [65], CViT [74], and MesoNet [11], but worse than the performance of MDS [15] and AVoiD-DF [72]. The superior performance of these methods may be due to the fact that they focus on the entire image frame rather than just the frontal face or lip regions.
Our study focuses on developing a multimodal deepfake detection method for frontal face videos. However, the DFDC test set contains various types of deepfake videos, including distant faces, side faces, multiple speakers, illumination variations, and low-light conditions. This variation poses challenges to the accurate extraction of faces and lips. Therefore, we took the frontal face videos from the DFDC test set and created a frontal face video test subset. As shown in Table XIII, our model achieves an AUC score of 83.84% on the frontal face video test subset. The results show that our proposed model remains highly effective for frontal face videos where facial identity or expression as well as lip movements may be manipulated, and shows reasonable performance on the entire test set. The results also show that our model has quite good generalization ability.
V Limitations
This study focuses on the authenticity detection of frontal face videos. Lack of frontal face, extreme environmental conditions, illumination variation, and multiple speakers in videos pose challenges to the face detection and lip capture modules in our method. For such types of videos, complete frame-based image analysis is required. However, in the proposed ensemble approach, the video-only model analyzes visual inconsistencies in facial movements, lighting, and other video content, independent of lip image data. While utilizing lip image data, the audio-visual model supplements the other two models by cross-referencing audio-visual synchronization to improve detection accuracy. Furthermore, our method is bimodal, requiring the input to contain both visual and acoustic modalities. To train our model, bimodal manipulations are required to train independent modality-specific networks and integrate their features into the decision-making module.
Due to the lack of suitable audio-visual datasets, it is currently not possible to test the generalization ability of the proposed model on more datasets. In addition to awaiting new publicly available audio-visual deepfake datasets, we are also preparing to expand the DeepfakeTIMIT dataset [76], where the audio in both real and fake videos is consistently authentic. We aim to apply various cutting-edge voice conversion technologies for acoustic manipulation.
As our ensemble approach consists of multiple transformers, the overall computational cost is higher compared to CNN-based or single transformer-based models in the literature. The number of trainable parameters in our video-only, audio-visual, and audio-only networks are 114.75M, 129.31M, and 87.73M respectively. The number of parameters in the linear layers of the decision-making module is 2.89M. In total, our proposed AVTENET model contains 334.68M trainable parameters. For inference time, we evaluate our proposed model on an NVIDIA GeForce RTX 2080 Ti GPU. To ensure accuracy, we executed the proposed model in the evaluation mode and performed multiple warm-up iterations before measurement to eliminate cold start overhead. The average inference time per video sample is 0.55 seconds. The recorded inference time only includes the forward pass of the proposed model and does not include any pre- or post-processing time.
VI Conclusions and Future Work
@@
In this study, we demonstrated the effectiveness of ensemble learning and studied deepfake video detection from an audio-visual perspective. We proposed a novel audio-visual transformer-based ensemble learning solution for effective and scalable video forgery detection. This approach is particularly effective for videos containing acoustic and visual manipulations, which cannot rely on existing methods because they can only detect unimodal manipulations or lack training diversity. Specifically, we devised a transformer-based model with a powerful detection ability due to its sequence modeling, parallelization, and attention mechanisms. We illustrated the performances of our multimodal transformer-based ensemble model on several manipulation techniques and compared them with various existing unimodal, multimodal, fusion, and ensemble models. The proposed model achieved state-of-the-art performance on the FakeAVCeleb dataset. To enhance trust and security in digital media environments, our model can be used to instantly detect deepfakes on social media platforms, which aids in content moderation and forensic analysis. Furthermore, the proposed solution can be extended through various foundation models trained using large amounts of data.
Acknowledgments
This work was supported in part by Academia Sinica under Grant AS-GC-111-M01, and by the National Council of Science and Technology, Taiwan, under Grants NSTC 113-2634-F-002-003 and NSTC 112-2221-E-001-009-MY3.